OpenStudy (anonymous):

Graph: (x-6)^2/4 - (y-4)^2/1=1 Please I have no idea what it would even look like :(

OpenStudy (anonymous):

can you write it out using the equation button?

OpenStudy (anonymous):

\[(x-6)^{2}\div4- (y-4)^{2}\div1=1\]

OpenStudy (anonymous):

Im supposed to graph it by hand :/

OpenStudy (anonymous):

can you combine the left side into 1 term?

OpenStudy (anonymous):

This is what it looks like.

OpenStudy (anonymous):

\[\frac{1}{4} (x-6)^2-(y-4)^2=1 \]Solve the above for y.\[y=\frac{1}{2} \left(8\pm \sqrt{32-12 x+x^2}\right) \]A plot is generated in two parts. The red part is the result of using the + sign and the blue the - sign. Calculations and Plot provided by Mathematica 8 Home Edition.

OpenStudy (anonymous):

THANK YOU SO MUCH :) What about Graph a hyperbola with foci at (0, ±10) and vertices at (0, ±8) using graphing technology ? :) think you can do it? :)

OpenStudy (anonymous):

Mathematica, and neither can I, process a statement like that directly. It is much easier and faster to plot equations using a computer program. The resulting curves are precise and accurate. There is a move afoot to introduce software like Mathematica into school math programs but it is not there yet by any stretch.